On January 1, 2019, ABC Company leased office equipment from ZZ, Inc. The lease terms require annual payments of $20,000 for 20 years with the first payment being due on December 31, 2019. The interest rate on the lease is 5%, and ABC will use the double-declining balance method to record the amortization of the leased asset. Assume the equipment had a 25 year remaining useful life at January 1, 2019 and the lease contract requires the equipment to be returned to ZZ, Inc. at the end of the lease. Calculate the amount of amortization expense recorded on the leased asset in 2020.
Answer #1

Answer: $ 22,432

PVA 5%, n=20 = [ { 1 - ( 1 / 1.05 ) 20 } / 0.05 ] = 12.46221

Present value of the lease payments = $ 20,000 x 12.4622 = $ 249,244

Rate of double declining balance amortization = 100 / 20 * 2 = 10 %

Amortization expense recorded on the leased asset in 2020 = $ 249,244 x 0.90 x 0.10 = $ 22,431.96